OpenStudy (anonymous):

Save My Work Write a paragraph proof that justifies the conditional statement “If 15 – 2x = 6 + x, then x =

OpenStudy (anonymous):

15-2x=6+x or, 15-2x-6 = 6+x -6 [Adding -6 on both sides] or, 9-2x = x or, 9 -2x+2x=x+2x [Adding 2x on both sides] or, 9 = 3X or 3x=9 [ Commutative rule] or 3x/3=9/3 [Dividing both sides by 3] or, x= 3 modify the reasons I gave in the brackets to make the paragraph.